# Asymptotic and compound e-values: multiple testing and empirical Bayes

### v3

@@ Abstract
We explicitly define the notions of (bona fide, approximate or asymptotic) compound p-values and e-values, which have been implicitly presented and extensively used in the recent multiple testing literature. While it is known that the e-BH procedure with compound e-values controls the FDR, we show the converse: every FDR controlling procedure can be recovered by instantiating the e-BH procedure with certain compound e-values. Since compound e-values are closed under averaging, this allows for combination and derandomization of FDR procedures. We then connect compound e-values to empirical Bayes. In particular, we use the fundamental theorem of compound decision theory to derive the log-optimal simple separable compound e-value for testing a set of point nulls against point alternatives: it is a ratio of mixture likelihoods. As one example, we construct asymptotic compound e-values for multiple t-tests, where the (nuisance) variances may be different across hypotheses. Our construction may be interpreted as a data-driven instantiation of the optimal discovery procedure (ODP), and our results provide the first type-I error guarantees for data-driven ODP.
@@ 1 Introduction
E-values provide an alternative inferential device to p-values. E-values encompass likelihood ratios and have provided some of the first non-trivial tests for composite null hypotheses, e.g., for testing if the data distribution is log-concave (Dunn et al., 2025); see Ramdas and Wang (2024) for a comprehensive introduction to e-values. In a multiple testing setting with hypotheses indexed by $\mathcal{K}=\{1,\ldots,K\}$ and null indices $\mathcal{N}\subseteq\mathcal{K}$, e-values $E_{1},\ldots,E_{K}$ can be used alongside the eBH (e-Benjamini-Hochberg) procedure of Wang and Ramdas (2022) to control the false discovery rate (FDR) of Benjamini and Hochberg (1995). The defining property of the e-values $E_{1},\ldots,E_{K}$ is that they are $[0,\infty]$-valued random variables satisfying $\mathbb{E}[E_{k}]\leq 1$ for all $k\in\mathcal{N}$. In this paper, we study $[0,\infty]$-valued random variables $E_{1},\ldots,E_{K}$ satisfying the compound constraint that
@@
holds exactly in finite samples, or satisfying suitable approximate or asymptotic versions of this constraint. We call $E_{1},\ldots,E_{K}$ (bona fide, approximate, or asymptotic) compound e-values. Compound e-values (without being given this name) have played an important role in the recent multiple testing literature. As one notable example, Ren and Barber (2024) solved the open problem of derandomizing the model-X knockoff filter (Candès et al., 2018) for variable selection with FDR control by reinterpreting the procedure in terms of compound e-values (see Section 4.3 for more details on derandomization).
@@
  • Compound e-values are central to multiple testing. In Section 4, we describe the fundamental role of compound e-values in multiple testing: Theorem 4.2 (in Section 4.1) shows how to construct compound e-values from any FDR-controlling procedure, while Theorem 4.5 (in Section 4.2) demonstrates that every FDR procedure can be recovered by e-BH applied to compound e-values. We explain how this framework enables combination and derandomization of multiple testing procedures (Section 4.3), including a novel application to derandomization of the randomized e-BH procedures of Xu and Ramdas (2024), and that compound e-values provide a natural notion of weights for the ep-BH procedure of Ignatiadis et al. (2024) (Section 4.4).
@@
[…]
  • Approximate and asymptotic compound e-values in multiple testing. While there are very few results on multiple testing with asymptotic and/or compound p-values, we show that the analogous notions with e-values are naturally compatible with guarantees for existing multiple testing procedures. Section 4 provides this unifying perspective. This is why we present Section 4 (on the central role of compound e-values in multiple testing) after Section 3 (which defines approximate and asymptotic compound e-values). In Section 7 we provide an alternative notion of approximate compound e-values that we call *approximate compound e-values. While these definitions coincide for single e-values, they differ in the compound case. The alternative definition is easier to verify but less compatible with existing multiple testing procedures.
@@
  • The simultaneous t-test sequence model. In Section 6 we demonstrate the usefulness of our definitions and constructions by instantiating them in the canonical setting of multiple t-tests where the nuisance parameters (variances) may be different across hypotheses. We first consider the normal setting (normal errors) with fixed sample-size for each hypothesis but growing number of hypotheses. Section 6.1.1 provides an explicit construction of CUI e-values, while Section 6.1.2 approximates the optimal discovery compound e-values using empirical Bayes, specifically, by pretending the variances are iid from a prior $G$ and estimating $G$ via the nonparametric maximum likelihood estimator (NPMLE) of Robbins (1950) and Kiefer and Wolfowitz (1956). Theorem 6.3 shows that this construction yields asymptotic compound e-values as $K\to\infty$. Our construction can be interpreted as a data-driven implementation of ODP, and to the best of our knowledge, Theorem 6.3 provides the first type-I error guarantee for data-driven ODP. The usefulness of these procedures is illustrated via a simulation study in Section 6.1.3 wherein we demonstrate substantial power-gains compared to alternatives such as UI and standard t-test e-values. In Section 6.2 we provide constructions that do not require normality: we provide a prototypical example of an asymptotic e-value (Section 6.2.1) that requires the sample size to grow to $\infty$, and then (Section 6.2.2) we provide an illustrative example of how to construct asymptotic compound e-values when either the per-hypothesis sample size or the number of hypotheses (or both) tends to $\infty$.
  • Connections to p-values. Throughout the paper, we establish connections between our definitions and results and their analogous counterparts for p-values. In Section 2 we define compound p-values, which have been previously called average significance controlling p-values by Armstrong (2022). Section 3 introduces approximate and asymptotic compound p-values, while Section 7 presents *approximate compound p-values; both of these concepts are new. A key finding is that the calibration between p-values and e-values established by Vovk and Wang (2021) extends to all variants introduced in this paper. We establish this correspondence across compound (Theorem 2.3), approximate (Theorem 3.3), asymptotic (Proposition 3.11), approximate compound (Proposition 3.14), and *approximate cases (Proposition 7.3). We also discuss results on multiple testing with compound p-values in Section 4.5, specifically generalizing a result of Armstrong (2022) on the Benjamini-Yekutieli procedure (BY) (Benjamini and Yekutieli, 2001) with compound p-values.
@@ 2 Compound p-values and e-values in multiple testing
The notions in Definition 2.1 are not new. The notion of compound p-variables was introduced by Armstrong (2022) under a different name (average significance controlling p-values and tests) and further studied in an empirical Bayes setting by Ignatiadis and Sen (2025). Meanwhile, the notion of compound e-variables was already used in the literature as a technical tool, for instance, it was first used (without being given a name) in Wang and Ramdas (2022, Theorem 3). The term compound e-values first appeared in Ignatiadis, Wang, and Ramdas (2024) with a distribution-dependent upper bound of $|\mathcal{N}(\mathbb{P})|$ instead of $K$ (typically the value of $|\mathcal{N}(\mathbb{P})|$ is not known except $|\mathcal{N}(\mathbb{P})|\leq K$). Other authors have used the term generalized e-values (Banerjee et al., 2023; Bashari et al., 2023; Zhao and Sun, 2024; Lee and Ren, 2024) or relaxed e-values (Ren and Barber, 2024; Gablenz and Sabatti, 2024).
@@
Suppose that $P_{k}$ is a p-variable for $\mathcal{P}_{k}$ for all $k\in\mathcal{K}$; $P_{1},\dots,P_{K}$ can be arbitrarily dependent. Moreover, let $E_{1},\dotsc,E_{K}$ be compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. If $(E_{1},\dots,E_{K})$ is independent of $(P_{1},\dots,P_{K})$, then $P_{1}/E_{1},\dots,P_{K}/E_{K}$ are compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. We revisit this example in Section 4.4.
@@
Suppose that $E_{1}^{(\ell)},\dots,E_{K}^{(\ell)}$ are compound e-variables for $\ell=1,\ldots,L$. Let $w_{1},\dotsc,w_{L}\geq 0$ be deterministic nonnegative numbers such that $\sum_{\ell=1}^{L}w_{\ell}=1$. Define $E_{k}=\sum_{\ell=1}^{L}w_{\ell}E_{k}^{(\ell)}$. Then $E_{1},\dots,E_{K}$ are compound e-variables.
@@ 3.2 Asymptotic p-values and e-values
We would also say “asymptotic e-values” (or p-values) in non-mathematical statements, and they can be interpreted as asymptotic e-variables (or p-variables) or their realizations. We sometimes simply say that a sequence is asymptotic for $\mathcal{P}$, and it should be clear from context whether we meant a sequence of asymptotic e-variables ($[0,\infty]$-valued) or that of asymptotic p-variables ($\mathbb{R}_{+}$-valued).111 Liu and Singh (1997) use the terminology “limiting p-values” for asymptotic p-values and “limiting p-values in the strong sense” for uniformly asymptotic p-values. Clearly, if $\mathcal{P}$ is finite, then asymptotic e-variables or p-variables are also uniformly asymptotic.
@@
We can naturally calibrate asymptotic p-variables and e-variables by applying a (possibly different) calibrator to each element in the sequence, in the sense of Theorem 3.3.
Proposition 3.11.
This result follows directly from Theorem 3.3.
The most common way of constructing asymptotic p-variables and e-variables is via limit theorems such as the central limit theorem.
Proposition 3.12.
@@
    $Z\sim|\mathrm{N}(0,1)|$, where we use the notation $|\mathrm{N}(0,1)|$ to denote the half-normal distribution. This would follow also from the central limit theorem and the continuous mapping theorem. [^2] An application of this case is as follows. In the setting of Proposition 3.12(ii), suppose $Z\sim\mathrm{N}(0,1)$ but we seek to use a function $h$ that is increasing in $|z|$ (rather than $z$) and that satisfies $\mathbb{E}[h(|Z|)]\leq 1$. Then the sequence $(E^{(n)})_{n\in\mathbb{N}}$ with $E^{(n)}=h(|Z^{(n)}|)$ is also a sequence of asymptotic e-variables.
@@
3.3 Approximate compound p/e-values
@@
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$ and $E_{1},\dots,E_{K}$ be $[0,\infty]$-valued random variables. The following two statements are equivalent (assuming that $\mathbb{P}$ is atomless for all $\mathbb{P}\in\mathcal{P}$):
  1. (E1)
    it holds that,
  2. (E2)
@@
3.4 Asymptotic compound p/e-values
@@ 4 Compound e-values are central to multiple testing
Our goal in this section is to demonstrate that compound e-values are fundamental to multiple testing and provide a useful conceptual framework for developing new multiple testing procedures that control the false discovery rate. To motivate what follows, we first record an observation about single hypothesis testing in which we seek to construct a level-$\alpha$ non-randomized test for a hypothesis $\mathcal{P}$, that is, a $\{0,1\}$-valued function $\phi(\cdot)$ such that $\mathbb{E}^{\mathbb{P}}[\phi(X)]\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}$. First, if $E=E(X)$ is an e-value for $\mathcal{P}$, then $\phi(X)=\mathds{1}_{\{E(X)\geq 1/\alpha\}}$ is a level-$\alpha$ test, since by Markov’s inequality $\mathbb{P}(E(X)\geq 1/\alpha)\leq\alpha\mathbb{E}^{\mathbb{P}}[E(X)]\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}$. Moreover, given any level-$\alpha$ test $\phi(X)$, we can construct an e-value via $E=\phi(X)/\alpha$ in which case $E\geq 1/\alpha$ if and only if $\phi(X)=1$. Thus, every test (including those based on thresholding p-values) implicitly uses an e-value and is recoverable by thresholding e-values at $1/\alpha$.
The remainder of this section develops a generalization of the above relation in the context of multiple testing. In Section 4.1 we show that compound e-values underlie every procedure controlling the false discovery rate (FDR), while in Section 4.2 we show that every FDR procedure can be recovered by applying the e-BH procedure of Wang and Ramdas (2022) to compound e-values. We then continue with practical applications of compound e-values to combination and derandomization of multiple testing procedures (Section 4.3) and explain why compound e-values provide the natural notion of weights for the weighted p-BH procedure (Section 4.4). We conclude by presenting a few known results on multiple testing with compound p-values (Section 4.5).
4.1 Compound e-values underlie every FDR procedure
@@
Benjamini and Hochberg (1995) proposed to control the false discovery rate, which is the expected value of the false discovery proportion, that is, $\mathrm{FDR}_{\mathcal{D}}^{\mathbb{P}}:=\mathbb{E}^{\mathbb{P}}[{F_{\mathcal{D}}}/{R_{\mathcal{D}}}]$ with the convention $/0=0$ 0.
@@
We make a few remarks. A notion related to approximate FDR control was used as proof technique (without being given that name) by Blanchard and Roquain (2008). A procedure with $(\varepsilon,\delta)$-approximate FDR control also satisfies
@@
The following result enables the construction of (approximate/asymptotic) compound e-values from any (approximate/asymptotic) FDR controlling procedure. The key observation underlying it has been made by several authors (Ren and Barber, 2024; Banerjee et al., 2023; Li and Zhang, 2023, 2025) for nonasymptotic FDR control.
Theorem 4.2 (From FDR control to compound e-values).
Let $\mathcal{D}$ be any procedure that controls the FDR at a known level $\alpha\in(0,1]$ for $(\mathcal{P}_{1},\dotsc,\mathcal{P}_{K})$ under $\mathcal{P}$. Define
$E_{k}=\frac{K}{\alpha}\frac{V_{k}}{R_{\mathcal{D}}\lor 1}.$ (4)
Then, $E_{1},\dots,E_{K}$ are compound e-variables for $(\mathcal{P}_{1},\dotsc,\mathcal{P}_{K})$ under $\mathcal{P}$. Further, if FDR control only holds $(\varepsilon,\delta)$-approximately or asymptotically, then $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)$-approximate or asymptotic compound e-values.
Proof.
Fix any $\mathbb{P}\in\mathcal{P}$. The result can then be seen from
where the inequality follows from the definition of FDR guarantee. The proof of the approximate and asymptotic claims follows analogously. ∎
4.2 Every FDR procedure is an instance of e-BH
@@
The following theorem due to Wang and Ramdas (2022) shows that e-BH controls the FDR.
@@
Notably, the above result requires no assumption whatsoever on the dependence across the $E_{k}$. The results in Wang and Ramdas (2022) handled $(\varepsilon,0)$-approximate compound e-values, so technically Theorem 4.4 offers a slight generalization in that respect.
Several authors including Ren and Barber (2024); Li and Zhang (2023, 2025) have observed that specific multiple testing procedures may be recast as e-BH with a specific choice of compound e-values as constructed in Theorem 4.2. In fact one can show, as we do next, that this holds for any multiple testing procedure that controls the FDR at a known level. This implies that if one seeks to control the FDR, it is without loss of generality to search among all procedures that apply e-BH to compound e-values.

### v4

@@ Abstract
We explicitly define the notions of (bona fide, approximate or asymptotic) compound p-values and e-values, which have been implicitly presented and used in the recent multiple testing literature. While it is known that the e-BH procedure with compound e-values controls the FDR, we show the converse: every FDR controlling procedure can be recovered by instantiating the e-BH procedure with certain compound e-values. Since compound e-values are closed under averaging, this allows for combination and derandomization of arbitrary FDR procedures. We then connect compound e-values to empirical Bayes. In particular, we use the fundamental theorem of compound decision theory to derive the log-optimal simple separable compound e-value for testing a set of point nulls against point alternatives: it is a ratio of mixture likelihoods. As one example, we construct asymptotic compound e-values for multiple t-tests, where the (nuisance) variances may be different across hypotheses. Our construction may be interpreted as a data-driven instantiation of the optimal discovery procedure, and our results provide the first type-I error guarantees for the same, along with significant power gains.
@@ 1 Introduction
E-values are an alternative inferential device to p-values. E-values encompass likelihood ratios and have provided some of the first non-trivial tests for composite null hypotheses, e.g., for testing if the data distribution is log-concave (Dunn et al., 2025); see Ramdas and Wang (2025) for a comprehensive introduction to e-values. In a multiple testing setting with hypotheses indexed by $\mathcal{K}=\{1,\ldots,K\}$ and null indices $\mathcal{N}\subseteq\mathcal{K}$, e-values $E_{1},\ldots,E_{K}$ can be used alongside the eBH (e-Benjamini-Hochberg) procedure of Wang and Ramdas (2022) to control the false discovery rate (FDR) of Benjamini and Hochberg (1995). The defining property of the e-values $E_{1},\ldots,E_{K}$ is that they are $[0,\infty]$-valued random variables satisfying $\mathbb{E}[E_{k}]\leq 1$ for all $k\in\mathcal{N}$. In this paper, we study $[0,\infty]$-valued random variables $E_{1},\ldots,E_{K}$ satisfying the compound constraint that
@@
holds exactly in finite samples, or satisfying suitable approximate or asymptotic versions of this constraint. We call $E_{1},\ldots,E_{K}$ (bona fide, approximate, or asymptotic) compound e-values.
Compound e-values (without being given this name) have played an important role in the recent multiple testing literature, after being introduced (almost in passing) in Wang and Ramdas (2022). As one notable example, Ren and Barber (2024) solved the open problem of derandomizing the model-X knockoff filter (Candès et al., 2018) for variable selection with FDR control by reinterpreting the procedure in terms of compound e-values (see Appendix C.1 for more details on derandomization).
@@
[…]
  • Compound e-values are central to multiple testing. In Section 4, we describe the fundamental role of compound e-values in multiple testing. Theorem 4.4 demonstrates the universality of e-BH applied to compound e-values: every FDR procedure can be recovered by applying e-BH to compound e-values. While there are very few results on multiple testing with asymptotic and/or compound p-values, we show that the analogous notions with e-values are naturally compatible with guarantees for existing multiple testing procedures. Section 4 provides this unifying perspective. This is why we present Section 4 (on the central role of compound e-values in multiple testing) after Section 3 (which defines approximate and asymptotic compound e-values). In Appendix C which accompanies Section 4, we explain how this framework enables combination and derandomization of multiple testing procedures (Appendix C.1), that compound e-values provide a natural notion of weights for the ep-BH procedure of Ignatiadis et al. (2024) (Appendix C.2), and that compound e-values can be used as weights for merging p-values (Appendix C.3).
@@
  • The simultaneous t-test sequence model. In Section 6 we demonstrate the usefulness of our definitions and constructions by instantiating them in the canonical setting of multiple t-tests where the nuisance parameters (variances) may be different across hypotheses. We first consider the normal setting (normal errors) with fixed sample-size for each hypothesis but growing number of hypotheses. Section 6.1.1 provides an explicit construction of CUI e-values, while Section 6.1.2 approximates the optimal discovery compound e-values using empirical Bayes, specifically, by pretending the variances are iid from a prior $G$ and estimating $G$ via the nonparametric maximum likelihood estimator (NPMLE) of Robbins (1950) and Kiefer and Wolfowitz (1956). Theorem 6.3 shows that this construction yields asymptotic compound e-values as $K\to\infty$. Our construction can be interpreted as a data-driven implementation of ODP, and to the best of our knowledge, Theorem 6.3 provides the first type-I error guarantee for data-driven ODP. The usefulness of these procedures is illustrated via a simulation study in Section 7 wherein we demonstrate substantial power-gains compared to alternatives such as UI and standard t-test e-values. Section 6.2 provides constructions that do not require normality: we provide a prototypical example of an asymptotic e-value (Section 6.2.1) that requires the sample size to grow to $\infty$, and then (Section 6.2.2) we provide an illustrative example of how to construct asymptotic compound e-values when either the per-hypothesis sample size or the number of hypotheses (or both) tends to $\infty$.
  • Connections to p-values. Throughout the paper, we establish connections between our definitions and results and their analogous counterparts for p-values. In Section 2 we define compound p-values, which have been previously called average significance controlling p-values by Armstrong (2022). Section 3 introduces approximate, asymptotic compound and *approximate compound p-values; these concepts are new. We find that the calibration between p-values and e-values established by Vovk and Wang (2021) extends to all variants introduced in this paper. We establish this correspondence across compound (Theorem 2.3), approximate (Theorem 3.3), asymptotic (Proposition 3.12), approximate compound (Proposition 3.14), and *approximate compound cases (Proposition B.2). We also discuss results on multiple testing with compound p-values in Appendix C.4, generalizing a result of Armstrong (2022) on the Benjamini-Yekutieli procedure (BY) (Benjamini and Yekutieli, 2001) with compound p-values.
@@ 2 Compound p-values and e-values in multiple testing
The notions in Definition 2.1 are not new. The notion of compound p-variables [^1] Habiger and Peña (2014) use the term compound p-values with a different meaning; see Appendix A. was introduced by Armstrong (2022) under a different name (average significance controlling p-values and tests) and further studied in an empirical Bayes setting by Ignatiadis and Sen (2025). Meanwhile, the notion of compound e-variables was already used in the literature as a technical tool, for instance, it was first used (without being given a name) in Wang and Ramdas (2022, Theorem 3). The term compound e-values first appeared in Ignatiadis, Wang, and Ramdas (2024) with a distribution-dependent upper bound of $|\mathcal{N}(\mathbb{P})|$ instead of $K$ (typically the value of $|\mathcal{N}(\mathbb{P})|$ is not known except $|\mathcal{N}(\mathbb{P})|\leq K$). Other authors have used the term generalized e-values (Banerjee et al., 2023; Bashari et al., 2023; Zhao and Sun, 2024; Lee and Ren, 2024) or relaxed e-values (Ren and Barber, 2024; Gablenz and Sabatti, 2024).
@@
Suppose that $P_{k}$ is a p-variable for $\mathcal{P}_{k}$ for all $k\in\mathcal{K}$; $P_{1},\dots,P_{K}$ can be arbitrarily dependent. Moreover, let $E_{1},\dotsc,E_{K}$ be compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. If $(E_{1},\dots,E_{K})$ is independent of $(P_{1},\dots,P_{K})$, then $P_{1}/E_{1},\dots,P_{K}/E_{K}$ are compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. We revisit this example in Appendix C.2.
@@
Suppose that $E_{1}^{(\ell)},\dots,E_{K}^{(\ell)}$ are compound e-variables for $\ell=1,\ldots,L$. Let $w_{1},\dotsc,w_{L}\geq 0$ be deterministic nonnegative numbers such that $\sum_{\ell=1}^{L}w_{\ell}=1$. Define $E_{k}=\sum_{\ell=1}^{L}w_{\ell}E_{k}^{(\ell)}$. Then $E_{1},\dots,E_{K}$ are compound e-variables. Also see Li and Zhang (2025, Proposition 5).
@@ 3.2 Asymptotic p-values and e-values
We would also say “asymptotic e-values” (or p-values) in non-mathematical statements, and they can be interpreted as asymptotic e-variables (or p-variables) or their realizations. We sometimes simply say that a sequence is asymptotic for $\mathcal{P}$, and it should be clear from context whether we meant a sequence of asymptotic e-variables ($[0,\infty]$-valued) or that of asymptotic p-variables ($\mathbb{R}_{+}$-valued).222 Liu and Singh (1997) use the terminology “limiting p-values” for asymptotic p-values and “limiting p-values in the strong sense” for uniformly asymptotic p-values. Clearly, if $\mathcal{P}$ is finite, then asymptotic e-variables or p-variables are also uniformly asymptotic.
@@
The most common way of constructing asymptotic p-variables and e-variables is via limit theorems such as the central limit theorem. The following example is a consequence of Propositions 3.7 and 3.8.
Example 3.11.
@@
    $Z\sim|\mathrm{N}(0,1)|$, where we use the notation $|\mathrm{N}(0,1)|$ to denote the half-normal distribution. This would follow also from the central limit theorem and the continuous mapping theorem. [^3] An application of this case is as follows. In the setting of Example 3.11(ii), suppose $Z\sim\mathrm{N}(0,1)$ but we seek to use a function $h$ that is increasing in $|z|$ (rather than $z$) and that satisfies $\mathbb{E}[h(|Z|)]\leq 1$. Then the sequence $(E^{(n)})_{n\in\mathbb{N}}$ with $E^{(n)}=h(|Z^{(n)}|)$ is also a sequence of asymptotic e-variables.
@@
We end this subsection by noting that we can naturally calibrate asymptotic p-variables and e-variables by applying a (possibly different) calibrator to each element in the sequence, in the sense of Theorem 3.3.
Proposition 3.12.
This result follows directly from Theorem 3.3.
3.3 Approximate compound p-values and e-values
@@
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$ and $E_{1},\dots,E_{K}$ be $[0,\infty]$-valued random variables. If each $\mathbb{P}\in\mathcal{P}$ is atomless, then Definition 3.13(ii) is equivalent to:
$\mathbb{E}^{\mathbb{P}}\left[\left(\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}E_{k}\right)\wedge t\right]\leq K(1+\varepsilon_{\mathbb{P}})+\delta_{\mathbb{P}}t\qquad\mbox{for all $t\in\mathbb{R}_{+}$ and all $\mathbb{P}\in\mathcal{P}$}.$
@@
We end this subsection by introducing an alternative definition in lieu of Definition 3.13 for approximate compound p-variables and e-variables. This definition has the advantage that it is easier to satisfy, but it provides weaker guarantees. We include it because the asymptotic notion it leads to for compound p-variables was central to the work of Ignatiadis and Sen (2025). However, we emphasize that in the remainder of the main text, we focus on Definition 3.13, since, as we will see in Section 4, approximate compound e-values as per Definition 3.13 are naturally compatible with approximate false discovery rate guarantees for the e-BH procedure.
  1. (i)
  2. (ii)
When all $\mathbb{P}\in\mathcal{P}$ are atomless and $K=1$, then Definitions 3.17 and 3.13 are equivalent to each other and equivalent to Definition 3.1 on approximate p-variables and e-variables. The definitions also agree when $\delta=0$, for instance, $(\varepsilon,0)^{*}$-approximate compound e-variables are identical to $(\varepsilon,0)$-approximate compound e-variables. Definition 3.17 is implied by Definition 3.13, as can be seen most clearly for approximate compound e-variables by comparing to the result of Proposition 3.15. In Appendix B we provide further results regarding Definition 3.17.
3.4 Asymptotic compound p-values and -values
@@ 4 Compound e-values are central to multiple testing
Our goal in this section is to demonstrate that compound e-values are fundamental to multiple testing and provide a useful conceptual framework for developing new multiple testing procedures that control the false discovery rate (FDR). To motivate what follows, we first record an observation about single hypothesis testing in which we seek to construct a level-$\alpha$ non-randomized test for a hypothesis $\mathcal{P}$, that is, a $\{0,1\}$-valued function $\phi(\cdot)$ such that $\mathbb{E}^{\mathbb{P}}[\phi(X)]\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}$. First, if $E=E(X)$ is an e-value for $\mathcal{P}$, then $\phi(X)=\mathds{1}_{\{E(X)\geq 1/\alpha\}}$ is a level-$\alpha$ test, since by Markov’s inequality $\mathbb{P}(E(X)\geq 1/\alpha)\leq\alpha\mathbb{E}^{\mathbb{P}}[E(X)]\leq\alpha$ for all $\mathbb{P}\in\mathcal{P}$. Moreover, given any level-$\alpha$ test $\phi(X)$, we can construct an e-value via $E=\phi(X)/\alpha$ in which case $E\geq 1/\alpha$ if and only if $\phi(X)=1$. Thus, every level-$\alpha$ test (including those based on thresholding p-values) implicitly uses an e-value and is recoverable by thresholding that e-value at $1/\alpha$ (Ramdas and Wang, 2025, Chapter 2).
The remainder of this section develops a generalization of the above relation in the context of multiple testing. Our primary result is that every FDR procedure can be recovered by applying the e-BH procedure of Wang and Ramdas (2022) to compound e-values.
@@
Benjamini and Hochberg (1995) proposed to control the FDR, which is the expected value of the false discovery proportion, that is, $\mathrm{FDR}_{\mathcal{D}}^{\mathbb{P}}:=\mathbb{E}^{\mathbb{P}}[{F_{\mathcal{D}}}/{R_{\mathcal{D}}}]$ with the convention $/0=0$ 0.
@@
We make a few remarks. A notion related to approximate FDR control was used as proof technique (without being given that name) by Blanchard and Roquain (2008).
A procedure with $(\varepsilon,\delta)$-approximate FDR control also satisfies
@@
The following theorem records the fact that e-BH controls the FDR.
@@
Notably, the above result requires no assumption whatsoever on the dependence across the $E_{k}$. The results in Wang and Ramdas (2022) handled $(\varepsilon,0)$-approximate compound e-values, so Theorem 4.3 offers a slight generalization in that respect.
Several authors including Ren and Barber (2024); Li and Zhang (2025) have observed that specific multiple testing procedures may be recast as e-BH with a specific choice of compound e-values. In fact one can show, as we do next, that this holds for any multiple testing procedure that controls the FDR at a known level. This implies that if one seeks to control the FDR, it is without loss of generality to search among all procedures that apply e-BH to compound e-values.